OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (–2, –3) and has a slope of 1/2? a) 2x+y=4 b) 2x-y=4 c) x+2y=4 d) x-2y=4

OpenStudy (lgbasallote):

you can use the formula \[y - y_1 = m(x-x_1)\]

OpenStudy (anonymous):

Use point-slope form \[y - y _{1} = m(x - x _{1}\] y + 3 = 1/2(x + 2) Standard form is Ax + By = C y + 3 = 1/2x + 1 Get it into standard form. 1/2x - y = 2 x - 2y = 4